Sixteen years ago Maddie was 50% younger than she is today. How old is she now?
Answer by jim_thompson5910(35256)

x - 16 = 0.5*x

-16 = 0.5x - x

-16 = -0.5x

x = -16/(-0.5)

x = 32

So today, she is 32 years old.